OpenStudy (anonymous):

I am stuck on trying to solve the following: 4/5 |4x+1| +1/2 = 3/5 A breakdown of the problem would be very helpful. Thanks!

OpenStudy (loser66):

\[\dfrac{4}{5|4x+1|}+\dfrac{1}{2}=\dfrac{3}{5}\] or \(\dfrac{4}{5}|4x+1|+\dfrac{1}{2}=\dfrac{3}{5}\)